# NACHOS: Neural Architecture Search for Hardware Constrained Early Exit Neural Networks

### v1

@@ V-A Dataset
The datasets used in the experimental comparisons are CIFAR-10 [49], CINIC-10 [50], and SVHN [51]. They consist of $32\times 32$ color images representing 10 classes of objects. Regarding CIFAR-10, the training set comprises 60000 images, while the testing set comprises 10000 images. Regarding CINIC-10, both the training set and the test set comprise 90000 images. Regarding SVHN, the training set comprises 73257 images, while the testing set comprises 26032 images.
@@
TABLE I: Results for the experimental campaign of NACHOS on CIFAR-10 dataset. Column labels explained in Table VI.
Model NAS B $F_{M}(\widetilde{x})$ (M) $F_{M}(x)$ $F_{A}(\widetilde{x})$ (%) $F_{A}(x)$ (%) $\theta$ U (%) Confidence
NACHOS yes 2 2.44 5.95 72.86 71.41 [0.2] [99.0, 1.0] Cumulative
  EDANAS [25] yes 5 2.46 5.99 67.76 67.34 [0.2,0.2,0.1,0.1] [0, 19.0, 61.0, 11.0, 9.0] Score margin
EECNN [52] no 2 7.48 25.14 76.50 77.20 [0.1] [100.0, 0.0] Max softmax
EEAlexnet [32] no 5 233.46 464.48 73.03 83.85 [0.7, 0.6, 0.6, 0.6] [38.0, 27.0, 19.0, 4.0, 12.0] Cumulative
EEResnet20 [32] no 10 21.25 315.26 69.71 79.91 [0.7, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6] [6.0, 0.0, 0.0, 33.0, 16.0, 18.0, 8.0, 9.0, 5.8, 4.2] Cumulative
TABLE II: Results for the experimental campaign of NACHOS on CINIC-10 dataset. Column labels explained in Table VI.
Model NAS B $F_{M}(\widetilde{x})$ (M) $F_{M}(x)$ $F_{A}(\widetilde{x})$ (%) $F_{A}(x)$ (%) $\theta$ U (%) Confidence
NACHOS yes 2 2.42 5.97 60.77 59.65 [0.2,0.1] [96.0, 4.0] Cumulative
  EDANAS [25] yes 4 2.43 6.06 60.52 59.52 [0.2,0.1,0.1] [36.0, 45.0, 9.5, 9.5] Score margin
EEAlexnet [32] no 5 300.20 464.48 64.35 73.30 [0.8, 0.8, 0.8, 0.8, 0.8] [16.0, 34.0, 13.0, 20.0, 17.0] Cumulative
EEResnet20 [32] no 10 142.57 315.26 61.69 69.01 [0.7, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6] [0.0, 0.0, 0.0, 0.0, 14.0, 24.0, 20.0, 0.4, 28.0, 10.0] Cumulative
TABLE III: Results for the experimental campaign of NACHOS on SVHN dataset. Column labels explained in Table VI.
Model NAS B $F_{M}(\widetilde{x})$ (M) $F_{M}(x)$ $F_{A}(\widetilde{x})$ (%) $F_{A}(x)$ (%) $\theta$ U (%) Confidence
NACHOS yes 3 1.49 4.61 79.98 88.71 [0.3,0.0,0.1] [56.0, 44.0, 0.0] Cumulative
  EDANAS [25] yes 5 1.48 6.01 78.10 86.10 [0.2, 0.1, 0.1, 0.1] [35.0, 57.0, 6.0, 1.0, 1.0] Score margin
EEAlexnet [32] no 5 263.51 464.48 89.00 93.10 [0.7, 0.6, 0.6, 0.6, 0.6] [23.0, 34.0, 27.0, 10.0, 6.0] Cumulative
EEResnet20 [32] no 10 70.20 315.26 88.53 93.28 [0.7, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6, 0.6] [0.0, 0.0, 0.0, 34.0, 13.0, 34.0, 11.0, 2.0, 2.0, 4.0] Cumulative
@@
TABLE IV: Computational costs for the experimental campaign of NACHOS on the three datasets.
Model CIFAR-10 CINIC-10 SVHN
NACHOS [2.41, 5.95] [2.28, 5.97] [1.17, 1.88, 4.61]
  EDANAS [25] [1.09, 1.44, 2.01, 4.02, 5.99] [1.51, 2.08, 4.09, 6.06] [1.11, 1.52, 2.04, 4.04, 6.01]
EEAlexnet [32] [87.27, 201.59, 380.82, 440.54, 464.48] Same Same
EEResnet20 [32] [19.57, 24.73, 29.89, 35.06, 43.99, 64.63, 85.28, 158.97, 239.48, 315.26] Same Same
TABLE V: ECE scores for the experimental campaign of NACHOS on CIFAR-10 dataset
@@ V-B Experimental setup
First, we report the most important terms used to describe the NACHOS framework in Table VI.
TABLE VI: List of the main symbols used in the experiments
Symbol Meaning
$\Omega_{x}$ Set of OFA backbones.
$\Omega_{\theta}$ Set of EEC hyperparameters.
$x$ Neural network architecture (backbone) sampled by $\Omega_{x}$.
$\theta$ The configuration of hyperparameters sampled by the EEC set.
$\widetilde{x}$ EENN sampled by $\Omega_{\widetilde{x}}$.
$F_{A}(\widetilde{x})$ The top1 accuracy of $\widetilde{x}$.
$F_{M}(\widetilde{x})$ The number of MACs of $\widetilde{x}$.
@@
  • the OFA supernet is based on MobileNetV3 and the hyperparameters of $\Omega_{x}$ refer to the depth, the kernel size, and the expansion rate while the resolution is fixed at 32 which is the size of the images of the datasets used in the experimental results;
  • the number of exit points (i.e., the points of the network where EECs are possibly placed) is set to 4;
  • the constraints are set as $\overline{F}_{A}=65\%$, $\overline{F}_{M}=2.7M$ on CIFAR-10; $\overline{F}_{A}=50\%,$ and $\overline{F}_{M}=2.7M$ on CINIC-10; $\overline{F}_{A}=75\%,$ and $\overline{F}_{M}=1.5M$ on SVHN. The constraints are chosen by considering the results of EDANAS on the different datasets.
@@
    the EENN found by NACHOS without imposing any constraint by selecting the best trade-off. We refer to this network as NACHOS_NC (NACHOS No Constraints).
@@
EDANAS has been run with the same configuration shown in  [25] but differently, the resolution of the networks has been fixed to $32\times 32$.
@@
The expert-designed networks have been trained with the following methodologies. EECNN has been trained by joint training. EEAlexNet and EEResnet20 have been trained by the same differentiable training adopted in NACHOS presented in the work [32]. In particular, the backbone is initially trained from scratch (i.e., no fine-tuning) for $\mu_{1}$ epochs, and then the standalone $L_{acc}$ loss function is optimized for $\mu_{2}+\mu_{3}$ epochs during the training of the whole EENN. The thresholds are chosen with the same grid search on the same intervals used in NACHOS but considering no constraints.
@@
We explain briefly the choice of the benchmarks. EDANAS is the only other framework in the literature optimizing jointly the backbone and the EECs. EECNN is a valid example of a CNN with a good trade-off between number of MACs and top1 accuracy. EEAlexNet and EEResnet20, trained for the same number of epochs of NACHOS (differently from EECNN), are taken to ensure fair benchmarking with state-of-the-art EENNs. Additionally, NACHOS is tested without $L_{peak}$ and $L_{cost}$ regularizations to verify their effectiveness in the framework.

### v2

@@ V-A Dataset
The datasets used in the experimental comparisons are CIFAR-10 [53], CINIC-10 [54], SVHN [55], and Imagenette [56], a subset of 10 easily classified classes from Imagenet [8], whose number of classes is 1000 and image resolution is 224. The first three datasets consist of $32\times 32$ color images while the resolution of Imagenette is $160\times 160$, and represent 10 classes of objects. Regarding CIFAR-10, the training set comprises 60000 images, while the testing set comprises 10000 images. Regarding CINIC-10, both the training set and the test set comprise 90000 images. Regarding SVHN, the training set comprises 73257 images, while the testing set comprises 26032 images. Regarding Imagenette, the training set comprises 9469 images while the testing set consists of 3925 images.
@@
TABLE II: Results for the experimental campaign of NACHOS on all the datasets. Column labels explained in Table I.
Dataset Model NAS B $F_{M}(\widetilde{x})$ (M) $F_{M}(x)$ (M) $F_{A}(\widetilde{x})$ (%) $F_{A}(x)$ (%) $\epsilon$ Confidence
SVHN NACHOS yes $3.00\pm 0.71$ $\bm{1.46\pm 0.03}$ $4.46\pm 0.18$ $79.96\pm 0.84$ $88.62\pm 0.23$ [$0.26\pm 0.05$, $0.13\pm 0.10$, $0.10$, –] Cumulative
EDANAS yes $4.60\pm 0.55$ $\underline{1.47\pm 0.02}$ $5.95\pm 0.05$ $77.98\pm 0.24$ $86.08\pm 0.19$ [$0.20\pm 0.00$, $0.14\pm 0.05$, $0.10\pm 0.00$, $0.10\pm 0.00$] Score margin
EEAlexnet no 5.00 263.51 464.48 $\bm{89.00}$ 93.10 [0.70, 0.60, 0.60, 0.60] Cumulative
EEResnet20 no 10.00 70.20 315.26 $\underline{88.53}$ 93.28 [0.70, 0.60, 0.60, 0.60, 0.60, 0.60, 0.60, 0.60, 0.60] Cumulative
CIFAR-10 NACHOS yes $2.80\pm 0.84$ $\bm{2.44\pm 0.02}$ $5.78\pm 0.19$ $72.65\pm 0.49$ $71.68\pm 0.31$ [$0.20\pm 0.00$, $0.24\pm 0.05$, $0.10\pm 0.00$, –] Cumulative
EDANAS yes $4.60\pm 0.55$ $\underline{2.47\pm 0.01}$ $6.00\pm 0.04$ $67.78\pm 0.12$ $67.48\pm 0.23$ [$0.20\pm 0.00$, $0.18\pm 0.04$, $0.12\pm 0.04$, $0.10\pm 0.00$] Score margin
EECNN no 2.00 7.48 25.14 $\bm{76.50}$ 77.20 [0.10] Max softmax
EEAlexnet no 5.00 233.46 464.48 $\underline{73.03}$ 83.85 [0.70, 0.60, 0.60, 0.60] Cumulative
EEResnet20 no 10.00 21.25 315.26 69.71 79.91 [0.7, 0.60, 0.60, 0.60, 0.60, 0.60, 0.60, 0.60, 0.60] Cumulative
CINIC-10 NACHOS yes $2.80\pm 0.84$ $2.42\pm 0.01$ $5.87\pm 0.10$ $60.85\pm 0.30$ $59.65\pm 0.11$ [$0.20\pm 0.04$, $0.22\pm 0.04$, $0.10\pm 0.00$, –] Cumulative
EDANAS yes $4.20\pm 0.45$ $2.45\pm 0.03$ $6.03\pm 0.02$ $60.65\pm 0.16$ $59.67\pm 0.17$ [$0.20\pm 0.00$, $0.18\pm 0.04$, $0.10\pm 0.00$, $0.10\pm 0.00$] Score margin
EEAlexnet no 5.00 300.20 464.48 $\bm{64.35}$ 73.30 [0.80, 0.80, 0.80, 0.80] Cumulative
EEResnet20 no 10.00 142.57 315.26 $\underline{61.69}$ 69.01 [0.70, 0.60, 0.60, 0.60, 0.60, 0.60, 0.60, 0.60, 0.60] Cumulative
Imagenette NACHOS yes $3.00\pm 0.71$ $\underline{92.34\pm 0.58}$ $104.09\pm 1.09$ $\underline{94.89\pm 0.15}$ $95.53\pm 0.08$ [–, $0.63\pm 0.06$, $0.68\pm 0.04$, $0.65\pm 0.07$] Cumulative
EDANAS yes $4.40\pm 0.55$ $\bm{81.92\pm 1.11}$ $231.34\pm 4.49$ $90.76\pm 0.54$ $96.53\pm 0.38$ [$0.10\pm 0.00$, $0.20\pm 0.00$, $0.14\pm 0.05$, $0.15\pm 0.07$] Score margin
EEEfficientNet-B0 no 4.00 182.77 201.13 $\bm{95.14}$ 97.25 [0.50, 0.50, 0.50] Cumulative
@@
TABLE III: Computational costs ($\gamma$) and utilization (U) for the experimental campaign of NACHOS on the four datasets.
Dataset Model Computational Costs $\gamma$ (M) Utilization U (%)
SVHN NACHOS [$1.12\pm 0.08$, $1.85\pm 0.11$, $2.5\pm 0.00$, –, $4.46\pm 0.18$] [$58.78\pm 21.89$, $39.91\pm 24.51$, $00.00\pm 0$ 0.00, –, $1.31\pm 2.92$]
EDANAS [$1.13\pm 0.12$, $1.51\pm 0.06$, $2.07\pm 0.04$, $4.01\pm 0.03$, $5.95\pm 0.05$] [$24.97\pm 16.69$, $82.94\pm 20.04$, $1.28\pm 2.64$, $0.24\pm 0.43$, $0.56\pm 0.55$]
EEAlexnet [87.27, 201.59, 380.82, 440.54, 464.48] [23.00, 34.00, 27.00, 10.00, 6.00]
EEResnet20 [19.57, 24.73, 29.89, 35.06, 43.99, 64.63, 85.28, 158.97, 239.48, 315.26] [0.00, 0.00, 0.00, 34.00, 13.00, 34.00, 11.00, 2.00, 2.00, 4.00]
EEEfficientNet-B0
CIFAR-10 NACHOS [$1.60\pm 0.0$, $2.18\pm 0.31$, $3.40\pm 0.44$, –, $5.78\pm 0.19$] [$41.31\pm 30.37$, $74.31\pm 30.32$, $00.00\pm 0$ 0.00, –, $0.90\pm 1.52$]
EDANAS [$1.11\pm 0.02$, $1.49\pm 0.04$, $2.03\pm 0.02$, $4.03\pm 0.03$, $6.00\pm 0.04$] [$0.87\pm 1.17$, $24.58\pm 7.56$, $51.06\pm 12.65$, $18.78\pm 10.17$, $5.08\pm 4.65$]
EEAlexnet [87.27, 201.59, 380.82, 440.54, 464.48] [38.00, 27.00, 19.00, 4.00, 12.00]
EEResnet20 [19.57, 24.73, 29.89, 35.06, 43.99, 64.63, 85.28, 158.97, 239.48, 315.26] [6.00, 0.00, 0.00, 33.00, 16.00, 18.00, 8.00, 9.00, 5.80, 4.20]
EEEfficientNet-B0 -
CINIC-10 NACHOS [$1.80\pm 0.20$, $2.45\pm 0.50$, $3.70\pm 0.42$, –, $5.87\pm 0.10$] [$69.75\pm 26.03$, $30.25\pm 26.03$, $00.00\pm 0$ 0.00, –, –]
EDANAS [$1.12\pm 0.00$, $1.48\pm 0.04$, $2.09\pm 0.03$, $4.06\pm 0.05$, $6.03\pm 0.02$] [$3.30\pm 0.00$, $25.74\pm 9.86$, $52.38\pm 6.46$, $15.8\pm 6.79$, $5.42\pm 3.48$]
EEAlexnet [87.27, 201.59, 380.82, 440.54, 464.48] [16.00, 34.00, 13.00, 20.00, 17.00]
EEResnet20 [19.57, 24.73, 29.89, 35.06, 43.99, 64.63, 85.28, 158.97, 239.48, 315.26] [0.00, 0.00, 0.00, 0.00, 14.00, 24.00, 20.00, 0.40, 28.00, 10.00]
EEEfficientNet-B0
Imagenette NACHOS [–, $46.00\pm 5.29$, $59.81\pm 3.63$, $70.21\pm 6.78$, $104.09\pm 1.09$] [$4.67\pm 3.33$, $14.31\pm 9.70$, $26.02\pm 14.98$, –, $72.42\pm 7.22$]
EDANAS [$59.58\pm 1.90$, $79.01\pm 1.83$, $92.86\pm 1.80$, $149.77\pm 4.38$, $231.34\pm 4.49$] [$0.37\pm 0.35$, $96.76\pm 1.86$, $0.70\pm 0.47$, $2.10\pm 0.42$, $1.48\pm 0.91$]
EEAlexnet
EEResnet20
EEEfficientNet-B0 [50.74, 103.16, 140.96, 201.13] [0.00, 6.34, 20.19, 73.47]
TABLE IV: ECE scores for the experimental campaign of NACHOS on CIFAR-10 dataset.
@@ V-B Experimental setup
  • the maximum number of iterations $N_{iter}$ is 30 and the number of epochs to trigger the convergence criterion $N_{tol}$ is 10;
  • considering our constrained mobile setting, we chose the pretrained MobileNetV3-based supernet, being the lightest in terms of MACs. We highlight that our framework is flexible enough to be applied to other CNN-based types of supernets. The hyperparameters of $\Omega_{x}$ refer to the depth, the kernel size, and the expansion rate (the architecture search space and building blocks are detailed in [37]) while the resolution is fixed at 32 which is the size of the images of the datasets used in the experimental results;
  • the number of exit points (i.e., the points of the network where EECs are possibly placed) is set to 4 since it is a reasonable arbitrary number of EECs in an EENN and is also chosen by EDANAS;
  • the constraints are set as $\overline{F}_{A}=65\%$, $\overline{F}_{M}=2.7M$ on CIFAR-10; $\overline{F}_{A}=50\%,$ and $\overline{F}_{M}=2.7M$ on CINIC-10; $\overline{F}_{A}=75\%,$ and $\overline{F}_{M}=1.5M$ on SVHN; $\overline{F}_{A}=90\%,$ and $\overline{F}_{M}=93M$ on Imagenette. The constraints are chosen by considering the results of EDANAS on the different datasets to test whether the searched models perform better in a similar setting.
  • the computational resources required: search time of 2 GPU days on a single GPU (NVIDIA A40) for both NACHOS and EDANAS. Their search times are the same since NACHOS takes more training time in the evaluation of the NNs but converges faster in terms of NAS epochs thanks to constraints as shown in Fig. 4.
  • the results of all the tested NAS methods, NACHOS and EDANAS, reports the mean and the standard deviation obtained by aggregating the results of the search over 5 different seeds following the best practice for NAS evaluations suggested in [57]. We ometted the mean and standard deviation of $\gamma$, $\epsilon$, and $U$ when the corresponding early exit has never been selected among the runs and placed a ’-’ instead.
Additionally, we report the hyperparameters of the genetic algorithm NSGA-II:
  • population size: 40;
  • initial population: non-dominated solutions of the previous NAS iteration;
  • crossover operator: two-point crossover with crossover probability=0.9;
  • mutation operator: polynomial mutation with distribution index=1.0;
  • selection operator: tournament selection with crowding distance sorting;
  • number of generations: 20;
  • survival: only the eight best solutions are kept according to ranking (Pareto dominance) and crowding distance sorting.
This setup allows effective convergence toward optimal trade-offs while preserving the diversity of the solutions, reducing the risk of premature convergence.
$N_{start},N_{iter}$, the hyperparameters of the genetic algorithm, and the training parameters such as batch size and learning rate follow the experimental setup of MSuNAS and EDANAS.
@@
Figure 3: Comparison between the Pareto fronts of runs of NACHOS enforcing and not enforcing constraints in the search on CIFAR-10 respectively, while setting the same seed.
Figure 4: Heatmap of the utilization of the second EEC of a model with three exits (B=3) found by NACHOS_NC varying the thresholds of the first two EECs.
@@
    the EENN found by NACHOS without using the $L_{cost}$ regularization by selecting the best trade-off. We refer to this network as NACHOS_NC (NACHOS No Constraints).
@@
  6. 6.
    EfficientNet-B0 [60] with 3 EECs placed by the computational-aware procedure explained in Algorithm 1. We refer to this network as EEEfficientNet-B0.
@@
EDANAS has been run with the same configuration shown in  [27] but differently, the resolution of the networks has been fixed to $160\times 160$ on Imagenette and $32\times 32$ on the other datasets.
@@
The expert-designed networks have been trained with the following methodologies. EECNN has been trained by joint training. EEAlexNet, EEResnet20, and EEEfficientNet-B0 have been trained by the same differentiable training adopted in NACHOS presented in the work [35]. In particular, the backbone is initially trained from scratch (i.e., no fine-tuning) for $\mu_{1}$ epochs, and then the standalone $L_{acc}$ loss function is optimized for $\mu_{2}+\mu_{3}$ epochs during the training of the whole EENN. The thresholds are chosen with the same grid search on the same intervals used in NACHOS but considering no constraints.
@@
We explain briefly the choice of the benchmarks. EDANAS is the only other framework in the literature optimizing jointly the backbone and the EECs. EECNN is a valid example of a CNN with a good trade-off between number of MACs and top1 accuracy. EEAlexNet, EEResnet20, and EEEfficientNet-B0 trained for the same number of epochs of NACHOS (differently from EECNN), are taken to ensure fair benchmarking with state-of-the-art EENNs. Additionally, NACHOS is tested without $L_{peak}$ and $L_{cost}$ regularizations to verify their effectiveness in the framework.
TABLE V: Results for the ablation studies of NACHOS on all datasets. Column labels are explained in Table I.
Dataset Model NAS B $F_{M}(\widetilde{x})$ (M) $F_{M}(x)$ (M) $F_{A}(\widetilde{x})$ (%) $F_{A}(x)$ (%) $\epsilon$ Confidence
SVHN NACHOS yes $3.00\pm 0.71$ $1.46\pm 0.03$ $4.46\pm 0.18$ $79.96\pm 0.84$ $88.62\pm 0.23$ [$0.26\pm 0.05$, $0.13\pm 0.10$, $0.10$, –] Cumulative
NACHOS_NC yes $2.80\pm 0.84$ $1.95\pm 0.10$ $4.74\pm 0.08$ $\bm{84.59\pm 0.47}$ $88.24\pm 0.27$ [$0.67\pm 0.06$, $0.16\pm 0.05$, $0.10\pm 0.00$, –] Cumulative
NACHOS_NP yes $2.80\pm 0.84$ $\bm{1.45\pm 0.02}$ $4.62\pm 0.18$ $76.94\pm 0.63$ $89.04\pm 0.43$ [$0.18\pm 0.08$, $0.07\pm 0.11$, $0.10\pm 0.00$, –] Cumulative
  CIFAR-10 NACHOS yes $2.80\pm 0.84$ $2.44\pm 0.02$ $5.78\pm 0.19$ $\bm{72.65\pm 0.49}$ $71.68\pm 0.31$ [$0.20\pm 0.00$, $0.24\pm 0.05$, $0.10\pm 0.00$, –] Cumulative
NACHOS_NC yes $3.00\pm 0.70$ $2.91\pm 0.11$ $6.43\pm 0.12$ $72.25\pm 0.22$ $71.38\pm 0.33$ [$0.65\pm 0.06$, $0.16\pm 0.05$] Cumulative
NACHOS_NP yes $2.40\pm 0.55$ $\bm{2.40\pm 0.05}$ $6.25\pm 0.42$ $71.38\pm 0.47$ $68.23\pm 0.59$ [–, $0.26\pm 0.05$, $0.15\pm 0.07$, –] Cumulative
  CINIC-10 NACHOS yes $2.80\pm 0.84$ $2.42\pm 0.01$ $5.87\pm 0.10$ $\bm{60.85\pm 0.30}$ $59.65\pm 0.11$ [$0.20\pm 0.04$, $0.22\pm 0.04$, $0.10\pm 0.00$, –] Cumulative
NACHOS_NC yes $3.00\pm 0.70$ $3.74\pm 0.13$ $6.85\pm 0.11$ $59.96\pm 0.12$ $59.51\pm 0.29$ [$0.70\pm 0.00$, $0.60\pm 0.17$, $0.58\pm 0.10$, –] Cumulative
NACHOS_NP yes $2.40\pm 0.55$ $\bm{2.41\pm 0.06}$ $4.79\pm 0.12$ $60.16\pm 0.34$ $59.63\pm 0.16$ [$0.20\pm 0.00$, $0.24\pm 0.05$, –, –] Cumulative
  Imagenette NACHOS yes $3.00\pm 0.71$ $92.34\pm 0.58$ $104.09\pm 1.09$ $94.89\pm 0.15$ $95.53\pm 0.08$ [–, $0.63\pm 0.06$, $0.68\pm 0.04$, $0.65\pm 0.07$] Cumulative
NACHOS_NC yes $3.20\pm 0.45$ $\bm{118.01\pm 1.47}$ $121.30\pm 3.15$ $\bm{97.36\pm 0.11}$ $96.46\pm 0.25$ [–, $0.70\pm 0.00$, $0.68\pm 0.04$, $0.66\pm 0.05$] Cumulative
NACHOS_NP yes $2.40\pm 0.55$ $84.38\pm 0.33$ $98.88\pm 0.41$ $85.24\pm 0.15$ $96.65\pm 0.25$ [–, $0.15\pm 0.06$, $0.17\pm 0.06$, –] Cumulative
TABLE VI: Computational costs ($\gamma$) and utilization (U) for the ablation studies of NACHOS on the four datasets.
Dataset Model Computational Costs $\gamma$ (M) Utilization U (%)
SVHN NACHOS [$1.12\pm 0.08$, $1.85\pm 0.11$, $2.53\pm 0.00$, –, $4.46\pm 0.18$] [$58.78\pm 21.89$, $39.91\pm 24.51$, $00.00\pm 0$ 0.00, –, $1.31\pm 2.92$]
NACHOS_NC [$1.16\pm 0.14$, $1.86\pm 0.18$, $2.60\pm 0.00$, –, $4.74\pm 0.08$] [$22.65\pm 2.61$, $72.29\pm 27.00$, $49.58\pm 0.00$, –, $4.20\pm 6.26$]
NACHOS_NP [$1.20\pm 0.07$, $2.03\pm 0.07$, $2.81\pm 0.00$, –, $4.62\pm 0.18$] [$78.29\pm 12.81$, $30.34\pm 6.92$, $00.00\pm 0$ 0.00, –, $3.50\pm 4.84$]
CIFAR-10 NACHOS [$1.60\pm 0.00$, $2.18\pm 0.31$, $3.40\pm 0.44$, –, $5.78\pm 0.19$] [$41.31\pm 30.37$, $74.31\pm 30.32$, $00.00\pm 0$ 0.00, –, $0.90\pm 1.52$]
NACHOS_NC [$1.68\pm 0.06$, $2.57\pm 0.38$, $3.67\pm 0.24$, –, $6.43\pm 0.12$] [$44.54\pm 22.45$, $42.02\pm 32.35$, $17.07\pm 2.23$, –, $15.41\pm 8.97$]
NACHOS_NP [–, $2.01\pm 0.27$, $3.05\pm 0.08$, –, $6.25\pm 0.42$] [$55.2\pm 1.91$, $74.39\pm 28.54$, –, –, $3.53\pm 3.27$]
CINIC-10 NACHOS [$1.80\pm 0.20$, $2.45\pm 0.50$, $3.70\pm 0.42$, –, $5.87\pm 0.10$] [$69.75\pm 26.03$, $30.25\pm 26.03$, $00.00\pm 0$ 0.00, –, –]
NACHOS_NC [$1.50\pm 0.00$, $3.03\pm 0.29$, $3.77\pm 0.12$, –, $6.85\pm 0.11$] [$9.87\pm 0.00$, $39.70\pm 27.63$, $74.61\pm 17.79$, –, $6.58\pm 11.93$]
NACHOS_NP [$1.25\pm 0.07$, $2.34\pm 0.12$, $2.95\pm 0.21$, –, $4.79\pm 0.12$] [$63.28\pm 2.86$, $60.85\pm 44.39$, $12.25\pm 0.95$, –, $8.94\pm 3.21$]
Imagenette NACHOS [–, $46.00\pm 5.29$, $59.81\pm 3.63$, $70.21\pm 6.78$, $104.09\pm 1.09$] [$4.67\pm 3.33$, $14.31\pm 9.70$, $26.02\pm 14.98$, –, $72.42\pm 7.22$]
NACHOS_NC [–, $52.35\pm 0.00$, $65.06\pm 4.30$, $92.06\pm 3.08$, $121.30\pm 3.15$] [–, $0.0\pm 0$ 0.0, $0.79\pm 1.09$, $8.77\pm 9.54$, $90.44\pm 8.88$]
NACHOS_NP [–, $51.00\pm 2.95$, $62.67\pm 2.52$, –, $98.88\pm 0.41$] [–, $24.55\pm 8.93$, $24.16\pm 12.65$, –, $65.86\pm 3.22$]